A Newly-acquired Equipment Requires An Annual Maintenance Costs Of P10,000. If The Annual Maintenance Cost Us Increased By 20% Each Year Every Year For 10 Years, What Is The Estimated Present Worth Of The Maintenance Costs If Money Is Worth 15%?


1
Expert's answer
2021-04-19T19:29:10-0400

Money worth "15\%" implies "15\%" is the opportunity cost of incurring maintenance cost. This is taken as the discount rate.

"PV=""10000\over 1.15""+" "10000\times 1.20\over 1.53^2" "+" "10000\times 1.20^2\over 1.15^3""+" "10000\times 1.20^3\over 1.15^4" "+" "...." "+" "10000\times 1.20^9\over 1.15^{10}"

"PV=" "10000\over 1.15" "(1+" "1.20\over 1.15" "+" "1.20^2\over 1.15^2" "+" "1.20^3\over1.15^3" "+" "....." "1.20^9\over 1.15^9" ")"

"PV=" "10000\over 1.15" "{(1-{{(1.20\over1.15)}}^{10}}\over{{1-{(1.20\over1.15)}}}" "=P106,100.509"